# Learning on LLM Output Signatures for Gray-Box Behavior Analysis

### v2

@@ Abstract
Large Language Models (LLMs) have achieved widespread adoption, yet our understanding of their behavior remains limited, particularly in detecting data contamination and hallucinations. While recently proposed probing techniques provide insights through activation analysis, they require “white-box” access to model internals, often unavailable. Current “gray-box” approaches typically analyze only the probability of the actual tokens in the sequence with simple task-specific heuristics. Importantly, these methods overlook the rich information contained in the full token distribution at each processing step. To address these limitations, we propose that gray-box analysis should leverage the complete observable output of LLMs, consisting of both the previously used token probabilities as well as the complete token distribution sequences - a unified data type we term LOS (LLM Output Signature). To this end, we develop a transformer-based approach to process LOS that theoretically guarantees approximation of existing techniques while enabling more nuanced analysis. Our approach achieves superior performance on hallucination and data contamination detection in gray-box settings, significantly outperforming existing baselines. Furthermore, it demonstrates strong transfer capabilities across datasets and LLMs, suggesting that LOS captures fundamental patterns in LLM behavior. Our code is available at: https://github.com/BarSGuy/LLM-Output-Signatures-Network.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse applications, yet their internal mechanisms remain poorly understood. This gap in understanding is particularly relevant in critical tasks like Hallucination Detection (HD) [45, 28, 17, 20, 40] and Data Contamination Detection (DCD) [6, 41, 58], where determining whether an LLM is fabricating information or has been exposed to specific training data is crucial for safe and reliable deployment.
@@
Figure 1: Left: The LLM processes the input “What does the cat chase?” and generates the output “A big mouse”. Right: The corresponding query/response (“resp.”) Token Distribution Sequences (TDS) and Actual Token Probabilities (ATP), together constituting the LLM Output Signature (LOS). We propose to analyze LLM behavior by learning directly from this unified representation.
@@
Previous work on LLM analysis has relied heavily on probing techniques that require restrictive white-box access to model internals [3, 35, 16, 15, 39]. Gray-box methods relax these assumptions by operating only on LLM outputs. Existing gray-box approaches typically analyze just the sequence of probabilities assigned to tokens that appear in the relevant input or output token sequence – a vector we term Actual Token Probabilities (ATP) [14, 22, 49, 18]. However, these methods, often based on heuristics, overlook the information contained in the complete Token Distribution Sequence (TDS) – a matrix holding the next-token probability distribution at each generation step, see Figure 1. This limitation can mask crucial differences in model behavior even at the level of a single time step. E.g., consider a model generating a token with probability 0.5 in two scenarios: in one case, the remaining next-token probability mass is concentrated on a single alternative $(0.5,0.5,0,...,0)$, while in the other it is spread across many tokens: $(0.5,0.01,...,0.01)$. These distributions suggest very different levels of model uncertainty, yet ATP-based approaches would treat them identically. Similarly, an ATP value of 0.1 at a certain time step could indicate either high uncertainty (if it is the highest probability in a diffused distribution) or strong evidence against the token (if it is a low-ranking probability in a peaked distribution). A recent promising approach [58] used some TDS information using heuristics, but a principled framework to utilize this data is still lacking.
@@
Our approach. We argue that a successful gray-box approach should leverage both ATP and TDS, together forming what we term the LLM Output Signature (LOS) (Figure 1) – the complete observable representation of LLM behavior in the gray-box setup. Instead of relying on heuristics, we treat LOS as a sequential, high-dimensional and structured data modality on which we apply principled deep learning techniques. We propose LOS-Net, a lightweight transformer encoder [^1] Around 1M parameters. operating on an effective encoding of ATP, TDS, and their interactions. We prove that LOS-Net can approximate a broad class of functions applied to the LOS of any LLM, subsuming many recent approaches [14, 22, 49, 18, 41, 58]. Our comprehensive empirical study on DCD and HD demonstrates a substantial information gap between using the complete LOS and relying solely on ATP. Notably, LOS-Net improves over all considered baselines across both tasks, often by a significant margin. Crucially, our architecture is extremely efficient, with detection times of $\approx 10^{-5}$s per instance, making it a compelling approach for applications such as on-line error detection for guided-generation. LOS-Net also exhibits promising dataset-level transfer and strong cross-model generalization, the latter suggesting its viable application to real-world tasks such as copyright-infringement detection over closed-source LLMs (see, e.g., our results on the BookMIA benchmark [41] in Section 5.2).
@@
Contributions are summarized as follows: (1) we introduce LOS as a suitable representation for analyzing LLM behavior, (2) we develop an effective and a learning framework for the LOS data modality, (3) we show this unifies and generalizes previous approaches, (4) we demonstrate it achieves superior performance across models, datasets, and tasks, and (5) exhibits strong empirical evidence for cross-model generalization and promising cross-dataset transfer abilities. The proposed LOS-Net proves effective for both HD and DCD, and its flexibility suggests broader potential for similar tasks while paving the way for foundational approaches to modeling LLM behaviors.
@@ 2 Related Work
We review background and related work on DCD and HD, focusing on studies leveraging logits or output probabilities. Given the breadth of research, we highlight the most relevant works for our setup and refer interested readers to Appendix C for further details on these tasks.
Data Contamination Detection. Early methods leveraged model loss [55, 8] for DCD, assuming that models overfit their training data. Later refinements introduced reference models—independent LLMs trained on disjoint datasets from a similar distribution—comparing their scores with the target model [9, 10]. However, this approach depends on the availability of a well-matched reference model (similar in its architecture), which is often impractical. Recently, [41] introduced Min-K%, which flags an input as contaminated if the log probability of its bottom K tokens exceeds a predefined threshold. Building on this approach, [58] proposed Min-K%++, which refines contamination detection by calibrating the next-token log-likelihood using the mean and standard deviation of log-likelihoods across all candidate tokens in the vocabulary.
Hallucination Detection has been studied as a means of enabling selective intervention, allowing LLMs to prevent fabricated outputs only when necessary [44, 56, 48]. Recently, [35] showed that training a classifier on top of LLMs’ hidden states is highly effective for hallucination detection. However, this method operates under the white-box assumption, requiring full access to the model’s internal components. In contrast, our paper explores a more constrained (gray-box) setting.
Output probability-Based Analysis. Previous works showed that using log probabilities or raw logits as decision thresholds can be effective for various tasks, including HD in LLMs [14, 49], correctness self-evaluation [22], uncertainty estimation [18], and zero shot learning [2]. However, these approaches often rely on naive handcrafted thresholding. Other approaches feed probabilities or logits into classifiers to get a more refined signal. Mosca et al. [34] computes logit differences for texts with and without a given word, training a classifier to detect adversarial attacks. Wu et al. [52] introduced LLMDet, which quantifies perplexity scores across models by analyzing next-token probabilities for selected n-grams, feeding these into a classifier to detect machine-generated content. Similarly, Verma et al. [51] presented Ghostbuster, which extracts token probabilities using simpler models and trains a linear classifier for the same aforementioned task. Both rely on linear classifiers and overlook the LLM’s TDS, limiting contextual understanding. In contrast, our method fully leverages textual context via the LOS for a more nuanced analysis.
@@ 3 Learning on LLM Output Signatures
3.1 Notation and Problem Formulation
Let $f$ denote a pretrained LLM, and $\vec{s}$ refer to a text input to $f$ consisting of $n$ tokens. When queried with $\vec{s}$, the LLM $f$ produces outputs $\mathbf{X}_{s}=f(\vec{s})$, i.e., a matrix in $\mathbb{R}^{n\times V}$ of next-token probabilities for each token in $\vec{s}$, where $V$ is the size of the token vocabulary. We define the LLM response to be $\vec{g}$ consisting of $m$ tokens generated using $f$’s outputs in $\mathbf{X}_{g}\in\mathbb{R}^{m\times V}$ (and $\mathbf{X}_{s}$). We refer to $\mathbf{X}_{s}$ or $\mathbf{X}_{g}$ as Token Distribution Sequences (TDS). See Figure 2. We also define $\mathbf{p}_{s}\in\mathbb{R}^{n},\mathbf{p}_{g}\in\mathbb{R}^{m}$, which holds the probabilities associated with the actual tokens appearing in $\vec{s},\vec{g}$ respectively. We denote these as the Actual Token Probabilities (ATP). Specifically, $(\mathbf{p}_{s})_{i}\coloneqq\mathbf{X}_{i,v}$ where $v\in\{1,\dots,V\}$ is the token used in the $i+1$ place in the sequence $\vec{s}$ and similarly for $\vec{g}$. See Figure 1 for an illustration. We call the pairs $(\mathbf{X}_{s},\mathbf{p}_{s})$ or $(\mathbf{X}_{g},\mathbf{p}_{g})$ the LLM Output Signature (LOS). For DCD, we analyze input sequences using $(\mathbf{X}_{s},\mathbf{p}_{s})$ since our interest lies in how the model processes the input text $\vec{s}$. In contrast, for HD, we use $(\mathbf{X}_{g},\mathbf{p}_{g})$ as we need to analyze the model’s generated response. We may use $(\mathbf{X},\mathbf{p})$ if the distinction between the tasks is irrelevant, and use $N$ as the sequence length.
@@ 3.1 Notation and Problem Formulation
3.2 Our Approach
Our approach consists of three main steps. Given an input $(\mathbf{X},\mathbf{p})$: (1) The probability distributions in TDS ($\mathbf{X}$) are sorted independently and sliced to only include the top $K$ ones at each time step, obtaining $\mathbf{X}^{\prime}$ ; (2) A learnable Rank Encoding $\text{RE}(\mathbf{X},\mathbf{p})$ is concatenated to $\mathbf{X}^{\prime}$ to capture relative probability information; (3) The resulting representation is processed by a lightweight transformer architecture, yielding the desired output. See illustration in Appendix E, Figure 9. In the remainder of this section, we provide a detailed explanation of each component.
Preprocessing the token distribution sequences.
Utilizing $\mathbf{X}$ may pose significant challenges due to three key factors. (1) Complexity: The vocabulary tensor can be extremely large in real-world scenarios. For instance, Liang et al. [27] (XLM-V) reported a vocabulary size of 1M tokens, which, for a small batch of documents and popular context sizes, would already entail processing a tensor of tens (or hundreds) of GBs. (2) Transferability: Vocabulary size and order may significantly vary between LLMs, something which can complicate transfer learning – e.g., training on one LLM and testing on another with a different vocabulary size; (3) Limited Access: In certain LLMs, such as those released by OpenAI, the output tensor $\mathbf{X}$ is only partially accessible, with APIs only exposing the (log-) probabilities for a small number of most likely tokens. To tackle these challenges, we propose selecting, for each row of $\mathbf{X}$, a fixed number of elements. Specifically, we preprocess $\mathbf{X}$ by sorting each row independently and selecting the top $K$ probabilities, as follows:
@@ 3.2 Our Approach
resulting in $\mathbf{X^{\prime}}\in\mathbb{R}^{N\times K}$. This approach not only reduces computational complexity but also provides a standardized representation that is independent of the vocabulary size (for an appropriate choice of $K$). Later, in Section 5, we will show how our approach can achieve strong empirical performance even for small values of $K$. Nevertheless, it is important to note that this preprocessing step removes the alignment of words across the vocabulary dimension. Exploring methods to retain or effectively utilize this alignment remains an avenue for future work.
Learnable Rank Encoding.
The tensor $\mathbf{X}^{\prime}$ provides a comprehensive description of the LLM’s output, but does not encode an important source of information: the probability $\mathbf{p}$ of the actual tokens appearing in the sequence, i.e, the ATP. The importance of this feature both in DCD and HD has already been demonstrated by a large body of prior work that operated only on this information [41, 14, 22, 49, 18]. Taking inspiration from these, we do also include ATPs as inputs to our architecture. However, we further complement these probabilities with additional information which allows us to contextualize them with respect to the whole TDS, i.e., $\mathbf{X}$. Specifically, we argue that valuable information is encoded in the rank (position) of the ATP within the vocabulary-wide (sorted) sequence of token probabilities. This information reveals both the model’s generation patterns and potential mismatches between predicted and actual tokens. The rank of the $i$-th token in the sequence is defined as: $r_{i}(\mathbf{X},\mathbf{p})=\sum_{v=1}^{V}\mathbb{I}(\mathbf{X}_{i,v}>p_{i}),$ where $\mathbb{I}(\cdot)$ is the indicator function. We encode the rank in a way to make this feature more amenable for learning, while still maintaining enough expressivity. Specifically, we first scale the rank between $[-1,1]$, obtaining $\mathbf{r}^{\text{scaled}}$. Then, we construct the following learnable rank encoding [^2] For certain DC datasets, we used a lookup table for Rank encoding, where the index corresponds to $r_{i}$ and the value is an embedding.,
Architecture.
Given the preprocessed TDS $\mathbf{X}^{\prime}$ and the developed rank encodings $\text{RE}(\mathbf{X},\mathbf{p})$, we first linearly project $\mathbf{X}^{\prime}$, concatenate it with $\text{RE}(\mathbf{X},\mathbf{p})$, and then feed it to an encoder-only transformer $\mathcal{T}$ with learnable positional encodings, operating in the temporal dimension [50]:

### v3

# Beyond Next Token Probabilities: Learnable, Fast Detection of Hallucinations and Data Contamination on LLM Output Distributions

@@ Abstract
The automated detection of hallucinations and training data contamination is pivotal to the safe deployment of Large Language Models (LLMs). These tasks are particularly challenging in settings where no access to model internals is available. Current approaches in this setup typically leverage only the probabilities of actual tokens in the text, relying on simple task-specific heuristics. Crucially, they overlook the information contained in the full sequence of next-token probability distributions. We propose to go beyond hand-crafted decision rules by learning directly from the complete observable output of LLMs — consisting not only of next-token probabilities, but also the full sequence of next-token distributions. We refer to this as the LLM Output Signature (LOS), and treat it as a reference data type for detecting hallucinations and data contamination. To that end, we introduce LOS-Net, a lightweight attention-based architecture trained on an efficient encoding of the LOS, which can provably approximate a broad class of existing techniques for both tasks. Empirically, LOS-Net achieves superior performance across diverse benchmarks and LLMs, while maintaining extremely low detection latency. Furthermore, it demonstrates promising transfer capabilities across datasets and LLMs. Full code is available at https://github.com/BarSGuy/Beyond-next-token-probabilities.
Guy Bar-Shalom* Fabrizio Frasca*
Technion Technion
guybs99@gmail.com fabriziof@campus.technion.ac.il
Derek Lim Yoav Gelberg Yftah Ziser Ran El-Yaniv Gal Chechik
MIT CSAIL Technion Nvidia Technion, Nvidia Bar-Ilan University, Nvidia
Haggai Maron
Technion, Nvidia
11footnotetext: Equal contribution.
@@ 1 Introduction
As the remarkable capabilities of LLMs continue to drive their expanding range of applications, detecting hallucinations [46, 30, 19, 22, 42], and training data contamination [7, 43, 58] becomes increasingly important to their reliable deployment and responsible use. Specifically, the tasks of Hallucination and Data Contamination Detection (resp. HD, DCD) relate to determining whether an LLM is fabricating information or is providing an incorrect answer to a user question, or whether it has been exposed to specific training data, such as copyrighted material.
@@
Figure 1: Left: The LLM processes the input “What does the cat chase?” and generates the output “A big mouse”. Right: The corresponding query/response Token Distribution Sequences (TDS) and Actual Token Probabilities (ATP), together constituting the LLM Output Signature (LOS). We propose to detect instances of hallucinations and data contamination by learning directly over this unified data representation, beyond task specific heuristics operating on partial information thereof.
@@
Prominent methods to tackle HD include probing techniques which, although effective, require restrictive white-box access to model internals [4, 37, 18, 17, 41], such as its hidden states. On both HD and DCD, gray-box methods relax these assumptions by operating only on LLM outputs, thus finding application to a broader set of models. These approaches [16, 24, 49, 20] typically extract simple features on the sequence of token probabilities, a vector we term Actual Token Probabilities (ATP). However, these methods, often based on heuristics, overlook the information contained in the complete next-token probability distributions generated over the token vocabulary at each generation step – we term this matrix the Token Distribution Sequence (TDS), see Figure˜1. Importantly, this limitation can mask distinctive patterns in the model’s text generation process, including its confidence or uncertainty, known to correlate with its correctness [25, 13]. This aspect is evident even at the level of a single generation step. Consider, e.g., an LLM generating a token with probability $0.5$ in two scenarios: in one case, the remaining next-token probability mass is concentrated on a single alternative (0.5,0.5,0,,0)(0.5,0.5,0,...,0), while in the other it is spread across many tokens: (0.5,0.01,,0.01)(0.5,0.01,...,0.01). See Figure˜12 in Appendix˜F for an illustration of a similar case. Yet ATP-based approaches would treat them identically. Similarly, an ATP value of $0.1$ at a certain time step could indicate either high uncertainty (if it is the highest probability in a diffused distribution) or strong evidence against the token (if it is a low-ranking probability in a peaked distribution). A recent promising approach [58] used some TDS information using heuristics, but a principled framework to utilize this data source is still lacking.
@@
Our approach.  We argue that a successful gray-box detection approach should leverage both ATP and TDS, together forming what we term the LLM Output Signature (LOS) (Figure˜1) – the complete observable representation of an LLM in the gray-box setup. Instead of relying on heuristics, we treat LOS as a sequential, high-dimensional and structured data modality on which we apply principled deep learning techniques. We propose LOS-Net, an efficient attention-based model [^1] Our model features around 1M parameters. operating on an effective encoding of ATP, TDS, and their interactions. We prove that LOS-Net can approximate a broad class of functions applied to the LOS, subsuming many recent approaches [16, 24, 49, 20, 43, 58]. Our comprehensive empirical study on DCD and HD demonstrates a substantial performance gap between using the complete LOS and relying solely on the ATP. Notably, LOS-Net improves over all considered baselines across both tasks, often by a significant margin. Crucially, our architecture is extremely efficient, with detection times of $\sim 10^{-5}$s per instance. This makes it a compelling approach for applications such as on-line error detection for guided-generation, as opposed to previously proposed popular methods based on multiple LLM prompting or generations, such as Semantic Entropy (SE) and P(True) [25, 24]. LOS-Net also exhibits promising dataset-level transfer and strong cross-model generalization, the latter suggesting its viable application to real-world tasks such as copyright-infringement detection over closed-source LLMs (see, e.g., our results on the BookMIA benchmark [43] in Section˜5.2). Last, we show LOS-Net retains strong performance also when processing a small subset of the TDS, expressed in terms of the number of top-scoring output probabilities at each generation step. This extends its successful application to LLMs with restricted API access, such as GPT models [36].
@@
Contributions.  (1) We introduce LOS as a suitable “data type” for the detection of hallucinations and data contamination, and develop LOS-Net, an effective and efficient learning framework for that. (2) We show this framework unifies and generalizes previous approaches, and demonstrate it achieves superior performance across models, datasets, and tasks. (3) We find that LOS-Net exhibits strong empirical evidence for cross-model generalization and promising cross-dataset transfer abilities.
@@ 2 Related Work
We review background and related work on DCD and HD, focusing on studies leveraging logits or output probabilities. Given the breadth of research, we highlight the most relevant works for our setup and refer interested readers toC for further details on these tasks.
Data Contamination Detection (DCD).  This task consists in identifying text passages an LLM has likely seen during training, or memorized. This is crucial for ensuring fair benchmarking of LLMs, guiding dataset curation, and auditing potential copyright infringement. Early methods leveraged model loss [55, 9], assuming that models overfit their training data. Later refinements introduced reference models – independent LLMs trained on disjoint datasets from a similar distribution – comparing their scores with the target model [10, 11]. However, this approach requires access to a well-matched reference model with similar architecture, which is often impractical in real-world settings. Recently, [43] introduced Min-K%, which flags an input as contaminated if the log probability of its bottom K tokens exceeds a predefined threshold. Building on this approach, [58] proposed Min-K%++, which refines contamination detection by calibrating the next-token log-likelihood using the mean and standard deviation of log-likelihoods across all candidate tokens in the vocabulary.
Hallucination Detection (HD).  This task has been studied to enable selective intervention, allowing LLMs to prevent fabricated outputs only when necessary [45, 56, 48]. Recently, [37, 3] showed that training a classifier on top of LLMs’ hidden states is highly effective for hallucination detection. However, these methods operate under the white-box assumption, requiring full access to the model’s internal states. In contrast, our paper explores a more constrained (gray-box) setting, of particular interest especially when targeting closed-source LLMs with restricted API access.
Output probability-Based Analysis.  Previous works showed that using log probabilities or raw logits as decision thresholds can be effective for various tasks, including HD in LLMs [16, 49], correctness self-evaluation [24], uncertainty estimation [20], and zero shot learning [2]. However, these approaches often rely on naive handcrafted thresholding. Other works [51, 35] rely on linear classifiers over features extracted from LLM outputs aiming at tackling adjacent tasks, such as detecting machine-generated text [52] but overlook the full TDS, limiting contextual understanding.
@@ 3 Learning on LLM Output Signatures
In this section, we define the LLM Output Signature (LOS) and introduce LOS-Net, a novel architecture specifically designed to process LOS.
3.1 LLM Output Signatures (LOS)
Let $f$ denote a pretrained LLM, and $\vec{s}$ a text input to $f$ consisting of $n$ tokens. When queried with $\vec{s}$, $f$ produces outputs $\mathbf{X}_{s}=f(\vec{s})$, i.e., a matrix in $\mathbb{R}^{n\times V}$ of next-token probabilities for each token in $\vec{s}$, where $V$ is the size of the token vocabulary. We define $\vec{g}$ to be the LLM response to $\vec{s}$, consisting of $m$ tokens generated using $f$’s outputs in $\mathbf{X}_{g}\in\mathbb{R}^{m\times V}$ (and $\mathbf{X}_{s}$). We refer to $\mathbf{X}_{s}$ or $\mathbf{X}_{g}$ as Token Distribution Sequences (TDS). See Figure˜11, Appendix˜E. We also define $\mathbf{p}_{s}\in\mathbb{R}^{n},\mathbf{p}_{g}\in\mathbb{R}^{m}$, vectors which holds the probabilities associated with the actual tokens appearing in $\vec{s},\vec{g}$ respectively. We denote these as the Actual Token Probabilities (ATP). Specifically, $(\mathbf{p}_{s})_{i}\coloneqq\mathbf{X}_{i,v}$ where $v\in\{1,\dots,V\}$ is the token used in the $i+1$ place in the sequence $\vec{s}$ and similarly for $\vec{g}$ (see Figure˜1). We call the pairs $(\mathbf{X}_{s},\mathbf{p}_{s})$ or $(\mathbf{X}_{g},\mathbf{p}_{g})$ the LLM Output Signature (LOS). For DCD, we analyze input sequences using $(\mathbf{X}_{s},\mathbf{p}_{s})$ as our interest lies in how the model processes the input text $\vec{s}$. For HD, we use $(\mathbf{X}_{g},\mathbf{p}_{g})$ as we need to make predictions on the model’s response. We may use $(\mathbf{X},\mathbf{p})$ if the distinction between the tasks is irrelevant, and use $N$ as the sequence length.
@@ 3.1 LLM Output Signatures (LOS)
3.2 LOS-Net
Learning from LOS data objects presents inherent challenges, particularly related to their encoding. Next, we detail these challenges and introduce our LOS-Net approach, illustrated in full in Appendix˜D and Figure˜10. What follows is a detailed explanation of each of its components.
Preprocessing the token distribution sequences.  Utilizing $\mathbf{X}$ may pose significant challenges due to three key factors. (1) Complexity: The vocabulary tensor can be extremely large in real-world scenarios. For instance, Liang et al. [29] (XLM-V) reported a vocabulary size of 1M tokens, which, for a small batch of documents and popular context sizes, would already entail processing a tensor of tens (or hundreds) of GBs. (2) Transferability: Vocabulary size and order may significantly vary between LLMs, something which can complicate transfer learning – e.g., training on one LLM and applying on another with a different vocabulary size; (3) Limited Access: As already mentioned, in certain LLMs, such as those released by OpenAI, the output tensor $\mathbf{X}$ is only partially accessible, with APIs only exposing a small number of the top (log-)probs. To tackle these challenges, we propose selecting, for each row of $\mathbf{X}$, a fixed number of elements. Specifically, we preprocess $\mathbf{X}$ by sorting each row independently and selecting the top $K$ probabilities, as follows:
@@ 3.2 LOS-Net
resulting in $\mathbf{X^{\prime}}\in\mathbb{R}^{N\times K}$. This approach not only reduces computational complexity but also provides a standardized representation independent of the vocabulary size (for an appropriate choice of $K$). Later, in Section˜5, we will show how even small values of $K$ can capture most of the TDS probability mass and enable strong empirical performance.
Encoding the ATP.  The tensor $\mathbf{X}^{\prime}$ provides a comprehensive description of the LLM’s output, but does not explicitly encode an important source of information: the probability $\mathbf{p}$ of the actual tokens appearing in the sequence, i.e, the ATP. While these values are technically present in the TDS (since it contains the full distribution), they are not directly distinguishable from the other token probabilities in the vocabulary. Thus, we do also include ATPs as separate inputs to our architecture and further complement these probabilities with additional information which allows us to contextualize them with respect to the whole TDS. Specifically, we argue that valuable information is encoded in the rank [^2] The rank of the $i$-th token is defined as: $r_{i}(\mathbf{X},\mathbf{p})=\sum_{v=1}^{V}\mathbb{I}(\mathbf{X}_{i,v}>p_{i}),$ where $\mathbb{I}(\cdot)$ is the indicator function. (position) of the ATP within the sorted TDS. This information reveals the “gap” between the actual token and the token the model would most likely expect to find instead. We encode the rank in a way to make this feature more amenable for learning: we apply scaling to a closed interval and transform it with specific parameters, obtaining $\text{RE}(\mathbf{X},\mathbf{p})$. More details are found in Appendix B.4.
Architecture.  Given the preprocessed TDS $\mathbf{X}^{\prime}$ and the rank encoding $\text{RE}(\mathbf{X},\mathbf{p})$, we first linearly project $\mathbf{X}^{\prime}$ via $\mathbf{W}\in\mathbb{R}^{K\times K^{\prime}}$, concatenate it with $\text{RE}(\mathbf{X},\mathbf{p})$, and then feed it to an encoder-only transformer module $\mathcal{T}$ with learnable positional encodings, operating in the sequence dimension [50]:
@@
Here, $\theta$ includes all model’s parameters, $\|$ denotes concatenation on the feature dimension. Finally, we pool over the [CLS] token and obtain output scores via a linear layer. The resulting model, LOS-Net, is trained with binary cross-entropy loss.
4 Generalization of Previous Approaches
As already mentioned, prior research has introduced various gray-box, methods for HD and DCD based on LLM’s output probabilities [16, 24, 49, 20]. In what follows, we propose a general framework to unify these diverse techniques, and show that this can be captured by our LOS-Net method, shedding light on its flexibility.
Motivating example: Min-K%  Shi et al. [43].  Min-K%, a prominent, recent method for DCD, makes predictions on an input text $\vec{s}$ based on a score $R$ calculated as the average of the smallest KK% log-probs: $R(\vec{s})=\frac{1}{|M|}\sum_{i\in M}\log(p_{i})$, with $M=\{i\mid p_{i}<\text{perc}(\mathbf{p},K)\}$ being the set of token indices whose probabilities are in the first $K$-th percentile of $\mathbf{p}$. We note that it is instructive to rewrite the scoring equation as: